An electric-dipole antenna, broadcasting microwave electromagnetic radiation for mobile communications, is located on the roof of a 111-m-high building and is oriented vertically. Two observers measure the signal strength (radiation intensity) from this antenna. Observer A is on the roof of another building, at the same height as the antenna and at a distance of 1.61 km from it. Observer B is on the street, with a direct-line distance of only 165 m from the antenna. What is the ratio of the signal strength measured by B, IB, to that measured by A, IA?
Answer #1

P_D = \frac{P_t}{4\pi r^2}

for A observer r_A = 1610m and for B observer r_B = \sqrt{(111m)^2 +(165m)^2 } = 198.86m so:

|I| = P_D

therefore:

\frac{|I_B|}{|I_A|} = \frac{r_A^2}{r_B^2} = \frac{(1610m)^2}{(198.86m)^2} = 65.54 \approx 66

\boxed{|I_B|= 66|I_A| }